For an acceptance sampling plan with n = 29 and c = 0, find the probability of accepting a lot that has a defect rate of 2%. (Round your answer to four decimal places.)

What is the probability of accepting the lot if the defect rate is 8%? (Round your answer to four decimal places.)
Answer #1

a)

Let X denote the number of defectives in the sample. Then X\sim Bin(n=29,p=0.02)

P(X=x)=\binom{29}{x}(0.02)^{x}(1-0.02)^{29-x},x=0,1,....,29

Required probability =

\\P(X\leq c)=P(X\leq 0)=P(X=0)=\binom{29}{0}(0.02)^{0}(1-0.02)^{29-0}\\ \\=0.5566

b)

Let X denote the number of defectives in the sample. Then X\sim Bin(n=29,p=0.08)

P(X=x)=\binom{29}{x}(0.08)^{x}(1-0.08)^{29-x},x=0,1,....,29

Required probability =

\\P(X\leq c)=P(X\leq 0)=P(X=0)=\binom{29}{0}(0.08)^{0}(1-0.08)^{29-0}\\ \\=0.0891